# The Decoupling of Binaries from Their Circumbinary Disks

### v2

@@ 1 Introduction
Nearly all massive galaxies are thought to harbor supermassive black holes (SMBHs) in their centers (e.g., Kormendy & Ho, 2013), resulting in the formation of SMBH binaries after the mergers of such galaxies and subsequent dynamical relaxation (e.g., Begelman et al., 1980; Mayer, 2013). The central regions of post-merger galaxies likely abound with dense gas (e.g. Barnes & Hernquist, 1992; Springel et al., 2005), which may form accretion disks around the binaries therein. Interactions between the binary and circumbinary disk may rapidly shrink the binary orbit (e.g., Haiman et al., 2009; Dittmann & Ryan, 2022), which is of particular importance when the binary is too tightly bound to inspiral efficiently due to gravitational interactions with field stars (Milosavljević & Merritt, 2003a, b) but too widely separated to inspiral efficiently due to gravitational wave (GW) emission.
@@
The evolution of accreting SMBH binaries is typically thought to span three phases: first, at early times and large binary separations, energy and angular momentum loss due to GW emission is negligible (Peters, 1964). During this stage, the orbital evolution of the binary is governed by interactions with the circumbinary disk (see, e.g., Lai & Muñoz, 2022, and references therein). Binary torques open a cavity in the circumbinary disk (e.g., Lubow, 1991; Miranda & Lai, 2015), the size of which is $\sim 1.75-3.25$ times larger than the binary semi-major axis depending on the binary mass ratio and disk viscosity (Ragusa et al., 2020; Dittmann & Ryan, 2022). On short timescales, depending on the angular momentum profile of the accreting gas, accretion onto the binary may be either suppressed (Artymowicz & Lubow, 1994; Ragusa et al., 2016) or enhanced, although this effect vanishes as the disk approaches a viscous steady state (Miranda et al., 2017; Dittmann & Ryan, 2022). During this first stage, the rate at which the binary inspirals is necessarily limited by that at which the disk viscously spreads.
@@
Eventually, the rate of change in the binary semi-major axis due to GW emission becomes sufficiently great that the binary and the disk decouple, and the orbital evolution of the binary is dominated by GWs. It has often been suggested that this stage occurs when the timescale for GW-driven inspiral ($\tau_{GW}$) becomes shorter than the viscous timescale ($\tau_{\nu}$) at the disk edge, and that this might lead to a dearth of gas during the SMBH merger, a cessation of emission at shorter wavelengths, and the interruption of jet activity (e.g., Liu et al., 2003; Milosavljević & Phinney, 2005; Farris et al., 2012). [^1] Although often cited when invoking a decoupling criterion of $\tau_{GW}\approx\tau_{\nu}$, Armitage & Natarajan (2002) actually suggested a velocity-based decoupling criterion. We more generally define decoupling as the epoch of disk-binary evolution during which the cavity size evolves in time more slowly than the binary semi-major axis. Regardless, eventually the inspiral proceeds quickly enough due to GW emission that the circumbinary disk is effectively frozen in place. Immediately after the binary merges, the mass lost from and kick imparted to the central black hole in may result in prompt electromagnetic counterparts (e.g. Schnittman & Krolik, 2008; O’Neill et al., 2009; Corrales et al., 2010), and subsequently as the circumbinary disk spreads the system will rebrighten as a single active galactic nucleus (e.g. Milosavljević & Phinney, 2005; Shapiro, 2010).
@@
Previous two-dimensional simulations have followed binaries through inspiral (e.g. Farris et al. (2015a) and Tang et al. (2018)), making pioneering predictions for variability during inspiral, but both used a single viscosity and assumed that the systems decouple when $\tau_{\rm GW}=\tau_{\nu}$.  [^2] Specifically, Farris et al. (2015a) implicitly rescaled the system parameters so that $\tau_{\rm GW}=\tau_{\nu}$ for their initial conditions. Rescaling their simulation to an initial separation of $100R_{g}$, their viscosity was intermediate between our choices of $\nu=0.01\,GM/c$ and $\nu=0.003\,GM/c$ (see Section 3), and their highest-resolution cells have $\Delta r\sim 2\,R_{g}$. Tang et al. (2018) used a viscosity roughly equivalent to our $\nu=0.01\,GM/c$ simulation, employed cells as small as $\Delta r\sim 0.375R_{g}$, and started with an initial separation of $30R_{g}$. See also recent work along the same vein, which also included the aforementioned post-merger effects (Krauth et al., 2023). Additionally, virtually all magnetohydrodynamic simulations of binary black holes, using an approximate but prescribed binary black hole metric or evolving the spacetime dynamically, have chosen initial conditions based on when the binary and disk are expected to decouple (e.g., Noble et al., 2012; Farris et al., 2012; Gold et al., 2014; Khan et al., 2018; Combi et al., 2022; Gutiérrez et al., 2022; Ruiz et al., 2023). Overestimating the binary separation at “decoupling” by a factor of $\sim 2-3$ may cause a mismatch between simulations and the physical systems that they intend to model. This inaccuracy can compound with differences between approximate viscous stresses and those following from magnetohydrodynamic turbulence that may also contribute to binaries decoupling later than predicted (e.g. Noble et al., 2012).
@@
An accurate understanding of when decoupling occurs and the subsequent system evolution is relevant to both present and future observational hunts for binary SMBHs (e.g. Graham et al., 2015; Charisi et al., 2016; Vaughan et al., 2016; Liu et al., 2018, 2019; Chen et al., 2020; Liao et al., 2021; Xin & Haiman, 2021). Furthermore, a more accurate understanding of the decoupling process will be useful for predicting and detecting optical counterparts to inspiraling binaries detected by LISA. Our results suggest electromagnetic counterparts to LISA events that may be sufficient to identify the host galaxies and redshifts thereof, making it possible to constrain the chirp mass of and luminosity distance to each GW source.
2 Analytical Arguments
@@
where $\Sigma$ is the surface density of the disk. In the following estimates we assume globally constant kinematic viscosity and surface density, noting both that some realistic disk models predict only weak dependence of $\nu$ on radius (e.g. Milosavljević & Phinney, 2005) and that the following reasoning holds for arbitrary kinematic viscosity profiles. Integrating the position of a fluid element from $r=r_{0}$ to $r=0$, we find the corresponding timescale to be
@@ 2.1 On Decoupling
A more appropriate criterion for decoupling is reached by equating $\dot{a}_{b}$ with $v_{r}(r_{c})$ (e.g. Armitage & Natarajan, 2002), resulting in the following approximation for the semi-major axis at which disk and binary decouple:
@@
We demonstrate in Section 4 that Equation (6) is significantly more accurate than Equation (5) when compared to our hydrodynamical simulations. Although a factor of $\sqrt{2\lambda\xi}$ is certainly within an order of magnitude in accuracy, taking $\xi\sim 2-3$ (as seen in simulations) and (conservatively) $\lambda=1$, this results in a factor of $\sim 2-3$ error in $a_{d}$ and a factor of $\sim 3-4$ error in the estimated orbital period at which the binary decouples.
2.2 On the Presence of Gas After Decoupling
@@
3 Numerical Methods
To quantitatively explore circumbinary disk decoupling, we carried out a suite of simulations equal mass binaries using the Athena++(Stone et al., 2020) code to solve the vertically-integrated Navier-Stokes equations in Cartesian coordinates. See Dittmann & Ryan (2023, in preparation) for a detailed explanation of our numerical methods. We initialized each binary with a separation of $a_{0}=100R_{g}$, and simulated disks with $\nu=\nu_{0}GM/c$, where $\nu_{0}\in\{0.1,0.03,0.01,0.003,0.001\}$. [^3] In terms of initial binary orbital parameters, $GM/c=a_{0}^{2}\Omega_{0}/10$, and thus our lowest viscosity was $\nu=10^{-4}\,{\rm a_{0}^{2}\Omega_{0}}$. In comparison to $\alpha-$viscosity models (Shakura & Sunyaev, 1973), which posit that $\nu=\alpha c_{s}H$ where $H$ is the disk scale height, or $\nu=\alpha(H/r)^{2}\sqrt{GMr}$ for a disk with constant $H/r$, $\nu_{0}\sim 10\alpha(H/r)^{2}$ at $r=a_{0}=100R_{g}$. Concretely, for disks with $H/r\sim 1/10,$ our $\nu_{0}=0.01$ model is roughly comparable to one with $\alpha=0.1$. In this parameterization, Equation (6) reduces to
@@ 3 Numerical Methods
We prescribed the motion of the binary analytically, using Equation (2) and the equations for the evolution of the binary angular velocity and orbital phase that follow therefrom. We use a softened gravitational potential for each black hole of $\Phi_{i}=-Gm(r_{i}^{2}+\epsilon_{g}^{2})^{-1/2},$ where $\epsilon_{g}={\rm max}(0.033a_{b},r_{g})$ is a gravitational softening length and $r_{i}$ is the distance between a fluid element and black hole $i$. We introduce a sink term for each black hole which removes gas at a rate $600(0.05a_{0}/r_{s})^{2}\nu_{0}\Omega_{0}(1-(r_{i}/r_{s})^{4})^{4}$ if $r_{i}<r_{s}$ and 00 otherwise, where the sink radius $r_{s}={\rm max}(0.05a_{b},6r_{g})$, $r_{g}=GM/2c^{2}$, and $\Omega_{0}=\sqrt{GM/a_{0}}$, so that at $r_{s}$ the mass removal timescale is a few times shorter than $\tau_{\nu}$. We use torque-controlled sinks (Dittmann & Ryan, 2021), removing specific spin angular momentum from accreting gas, with respect to each SMBH, equal to that at the innermost stable circular orbit ($6r_{g}$) if $r_{i}>6r_{g}$ and not altering fluid velocities if $r_{i}\leq 6r_{g}$. We implemented damping zones at the outer boundary and a locally isothermal equation of state similarly to Moody et al. (2019), setting the azimuthal Mach number to $\mathcal{M}=10$. We used the Roe approximate Riemann solver (Roe, 1981), third-order reconstruction (Felker & Stone, 2018), and the second-order van Leer (VL2) time integrator (Stone & Gardiner, 2009). We utilize the adaptive mesh refinement infrastructure within Athena++ to increase resolution as the binary contracts, maintaining a cell aspect ratio of $\Delta x/r=\Delta y/r\approx(2/300){\rm{max}}(a_{b},r)$ when possible as long as $r>{\rm max}(a_{b},20r_{g}).$
We initialized each disk with a surface density profile of $\Sigma=\Sigma_{0}\exp[-(r/2.5a_{0})^{-12}][1-0.7\sqrt{(}a_{0}/r)]$ (see, e.g. Lynden-Bell & Pringle, 1974; Miranda et al., 2017; Dittmann & Ryan, 2022), a radial velocity of $-3\nu/2r$, and an angular velocity profile of
@@ 4 Results
For the sake of brevity we do not report here a detailed analysis of gas torques on the binary during the inspiral. However, over the course of decoupling gravitational torques between the binary, circumbinary disk, and streams of gas within the cavity, which are largely responsible for negative contributions to the gravitational torque on the binary from the gas (e.g., Tiede et al., 2020; Dittmann & Ryan, 2022), become weaker as the binary decouples. The minidisks and their positive torque contributions remain, so the gravitational torque on the binary tends to increase by a factor of order unity (normalized by the accretion rate and binary angular momentum $\dot{M}J_{b}/M$) during the decoupling process, although this effect may be more substantial in thinner disks.
4.1 Disk Morphology
@@
We measured the evolution of the circumbinary disk cavity using time-averaged histograms of the fluid surface density, binned according to fluid element semi-major axis (see, e.g. Figure 17 of Dittmann & Ryan, 2022). We defined the semi-major axis of the cavity ($a_{\rm cav}$) as that at which surface density reaches a value of $\Sigma_{0}/5$ at $a>a_{b}$. Although our definition of cavity semi-major axis is somewhat arbitrary, we have verified that our results are insensitive to the precise definition used. The top panel of Figure 2 presents the results of these measurements for each viscosity, along with nonlinear least-squares fits of a smoothly broken power law shown using dashed lines. For our broken power law fit, we fix early- and late-time power law indices $a_{\rm cav}\propto a_{b}^{p}$ to $p=1$ and $p=0$ respectively, fitting for a break location, transition width, and constant of proportionality. We define the binary semi-major axis which characterizes the decoupling of the binary from the circumbinary disk ($a_{d}$) as the location of the power-law break. Although $a_{d}$ is a characteristic point in the evolution of the binary, Figure 2 also illustrates that decoupling is a gradual process as opposed to an instantaneous one.
The bottom panel of Figure 2 compares our measurements with the predictions made by Equations (5) and (6) using $\lambda=1$ and measuring $\xi$ at the end of the burn-in stage of each simulation. As expected based on Section 2.1, timescale-based reasoning overestimates the characteristic binary semi-major axis at which decoupling occurs. Additionally, Equation (6) makes fewer and weaker assumptions than Equation (5), not depending on $\lambda$ and depending more weakly on $\xi$. Comparing the values of $a_{d}$ in Figure 2 to the surface density snapshots in Figure 1 illustrates that the disk-binary system continues to evolve together after this nominal decoupling point ($a_{d}\approx 56\,R_{g}$ for the $\nu_{0}=0.003$ simulation), albeit at a slower rate.
4.2 Accretion
@@
In the $\nu_{0}=0.001$ case, periodicity is dominated by the orbital period of the disk cavity and its harmonics as has been observed previously (e.g. Farris et al., 2014; Noble et al., 2021). Although in the lab frame the cavity orbital frequency slightly increases between the beginning of the simulation and decoupling, as the cavity shrinks, the frequency relative to the binary decreases, and the binary evolution outpaces that of the disk. At late times, almost all lab-frame variability is dwarfed by the secular decline in the accretion rate. Our averaging window was also large enough, in terms of the binary orbital period, to identify some low-frequency periodicity due to accretion from the cavity walls. However, even in terms of the evolving binary period, there is virtually no variability on the orbital period of the binary. In these cases, however, effects such as Doppler modulation (e.g., D’Orazio & Haiman, 2017) and self-lensing (e.g., Davelaar & Haiman, 2022) might lead to variability in observed light curves.
4.3 Gravitational Waves

### v3

@@ I Introduction
Nearly all massive galaxies are thought to harbor supermassive black holes (SMBHs) in their centers (e.g., Kormendy and Ho, 2013), resulting in the formation of SMBH binaries after the mergers of such galaxies and subsequent dynamical relaxation (e.g., Begelman et al., 1980; Mayer, 2013). The central regions of post-merger galaxies likely abound with dense gas (e.g. Barnes and Hernquist, 1992; Springel et al., 2005), which may form accretion disks around the binaries therein. Interactions between the binary and circumbinary disk may rapidly shrink the binary orbit (e.g., Haiman et al., 2009; Dittmann and Ryan, 2022), which is of particular importance when the binary is too tightly bound to inspiral efficiently due to gravitational interactions with field stars (Milosavljević and Merritt, 2003a, b) but too widely separated to inspiral efficiently due to gravitational wave (GW) emission.
@@
The evolution of accreting SMBH binaries is typically thought to span three phases: first, at early times and large binary separations, energy and angular momentum loss due to GW emission is negligible (Peters, 1964). During this stage, the orbital evolution of the binary is governed by interactions with the circumbinary disk (see, e.g., Lai and Muñoz, 2022, and references therein). Binary torques open a cavity in the circumbinary disk (e.g., Lubow, 1991; Miranda and Lai, 2015), the size of which is $\sim 1.75-3.25$ times larger than the binary semi-major axis depending on the binary mass ratio and disk viscosity (Ragusa et al., 2020; Dittmann and Ryan, 2022). On short timescales, depending on the angular momentum profile of the accreting gas, accretion onto the binary may be either suppressed (Artymowicz and Lubow, 1994; Ragusa et al., 2016) or enhanced, although this effect vanishes as the disk approaches a viscous steady state (Miranda et al., 2017; Dittmann and Ryan, 2022). During this first stage, the rate at which the binary inspirals is necessarily limited by that at which the disk viscously spreads.
@@
Eventually, the rate of change in the binary semi-major axis due to GW emission becomes sufficiently great that the binary and the disk decouple, and the orbital evolution of the binary is dominated by GWs. It has often been suggested that this stage occurs when the timescale for GW-driven inspiral ($\tau_{GW}$) becomes shorter than the viscous timescale ($\tau_{\nu}$) at the disk edge, and that this might lead to a dearth of gas during the SMBH merger, a cessation of emission at shorter wavelengths, and the interruption of jet activity (e.g., Liu et al., 2003; Milosavljević and Phinney, 2005; Farris et al., 2012). [^1] Although often cited when invoking a decoupling criterion of $\tau_{GW}\approx\tau_{\nu}$, Armitage and Natarajan (2002) actually suggested a velocity-based decoupling criterion. We more generally define decoupling as the epoch of disk-binary evolution during which the cavity size evolves in time more slowly than the binary semi-major axis. Regardless, eventually the inspiral proceeds quickly enough due to GW emission that the circumbinary disk is effectively frozen in place. Immediately after the binary merges, the mass lost from and kick imparted to the central black hole in may result in prompt electromagnetic counterparts (e.g. Schnittman and Krolik, 2008; O’Neill et al., 2009; Corrales et al., 2010), and subsequently as the circumbinary disk spreads the system will rebrighten as a single active galactic nucleus (e.g. Milosavljević and Phinney, 2005; Shapiro, 2010).
@@
Previous two-dimensional simulations have followed binaries through inspiral (e.g. Farris et al. (2015a) and Tang et al. (2018)), making pioneering predictions for variability during inspiral, but both used a single viscosity and assumed that the systems decouple when $\tau_{\rm GW}=\tau_{\nu}$.  [^2] Specifically, Farris et al. (2015a) implicitly rescaled the system parameters so that $\tau_{\rm GW}=\tau_{\nu}$ for their initial conditions. Rescaling their simulation to an initial separation of $100R_{g}$, their viscosity was intermediate between our choices of $\nu=0.01\,GM/c$ and $\nu=0.003\,GM/c$ (see Section III), and their highest-resolution cells have $\Delta r\sim 2\,R_{g}$. Tang et al. (2018) used a viscosity roughly equivalent to our $\nu=0.01\,GM/c$ simulation, employed cells as small as $\Delta r\sim 0.375R_{g}$, and started with an initial separation of $30R_{g}$. See also recent work along the same vein, which also included the aforementioned post-merger effects (Krauth et al., 2023). Additionally, virtually all magnetohydrodynamic simulations of binary black holes, using an approximate but prescribed binary black hole metric or evolving the spacetime dynamically, have chosen initial conditions based on when the binary and disk are expected to decouple (e.g., Noble et al., 2012; Farris et al., 2012; Gold et al., 2014; Khan et al., 2018; Combi et al., 2022; Gutiérrez et al., 2022; Ruiz et al., 2023). Overestimating the binary separation at “decoupling” by a factor of $\sim 2-3$ may cause a mismatch between simulations and the physical systems that they intend to model. This inaccuracy can compound with differences between approximate viscous stresses and those following from magnetohydrodynamic turbulence that may also contribute to binaries decoupling later than predicted (e.g. Noble et al., 2012).
@@
An accurate understanding of when decoupling occurs and the subsequent system evolution is relevant to both present and future observational hunts for binary SMBHs (e.g. Graham et al., 2015; Charisi et al., 2016; Vaughan et al., 2016; Liu et al., 2018, 2019; Chen et al., 2020; Liao et al., 2021; Xin and Haiman, 2021). Furthermore, a more accurate understanding of the decoupling process will be useful for predicting and detecting optical counterparts to inspiraling binaries detected by LISA. Our results suggest electromagnetic counterparts to LISA events that may be sufficient to identify the host galaxies and redshifts thereof, making it possible to constrain the chirp mass of and luminosity distance to each GW source.
II Analytical Arguments
@@
where $\Sigma$ is the surface density of the disk. In the following estimates we assume globally constant kinematic viscosity and surface density, noting both that some realistic disk models predict only weak dependence of $\nu$ on radius (e.g. Milosavljević and Phinney, 2005) and that the following reasoning holds for arbitrary kinematic viscosity profiles. Integrating the position of a fluid element from $r=r_{0}$ to $r=0$, we find the corresponding timescale to be
@@ II.1 On Decoupling
A more appropriate criterion for decoupling is reached by equating $\dot{a}_{b}$ with $v_{r}(r_{c})$ (e.g. Armitage and Natarajan, 2002), resulting in the following approximation for the semi-major axis at which disk and binary decouple:
@@
We demonstrate in Section IV that Equation (6) is significantly more accurate than Equation (5) when compared to our hydrodynamical simulations. Although a factor of $\sqrt{2\lambda\xi}$ is certainly within an order of magnitude in accuracy, taking $\xi\sim 2-3$ (as seen in simulations) and (conservatively) $\lambda=1$, this results in a factor of $\sim 2-3$ error in $a_{d}$ and a factor of $\sim 3-4$ error in the estimated orbital period at which the binary decouples.
II.2 On the Presence of Gas After Decoupling
@@
III Numerical Methods
[…]
@@ III Numerical Methods
We prescribed the motion of the binary analytically, using Equation (2) and the equations for the evolution of the binary angular velocity and orbital phase that follow therefrom. We use a softened gravitational potential for each black hole of $\Phi_{i}=-Gm(r_{i}^{2}+\epsilon_{g}^{2})^{-1/2},$ where $\epsilon_{g}={\rm max}(0.033a_{b},r_{g})$ is a gravitational softening length and $r_{i}$ is the distance between a fluid element and black hole $i$. We introduce a sink term for each black hole which removes gas at a rate $600(0.05a_{0}/r_{s})^{2}\nu_{0}\Omega_{0}(1-(r_{i}/r_{s})^{4})^{4}$ if $r_{i}<r_{s}$ and 0 otherwise, where the sink radius $r_{s}={\rm max}(0.05a_{b},6r_{g})$, $r_{g}=GM/2c^{2}$, and $\Omega_{0}=\sqrt{GM/a_{0}}$, so that at $r_{s}$ the mass removal timescale is a few times shorter than $\tau_{\nu}$. We use torque-controlled sinks (Dittmann and Ryan, 2021), removing specific spin angular momentum from accreting gas, with respect to each SMBH, equal to that at the innermost stable circular orbit ($6r_{g}$) if $r_{i}>6r_{g}$ and not altering fluid velocities if $r_{i}\leq 6r_{g}$. We implemented damping zones at the outer boundary and a locally isothermal equation of state similarly to Moody et al. (2019), setting the azimuthal Mach number to $\mathcal{M}=10$. We used the Roe approximate Riemann solver (Roe, 1981), third-order reconstruction (Felker and Stone, 2018), and the second-order van Leer (VL2) time integrator (Stone and Gardiner, 2009). We utilize the adaptive mesh refinement infrastructure within Athena++ to increase resolution as the binary contracts, maintaining a cell aspect ratio of $\Delta x/r=\Delta y/r\approx(2/300){\rm{max}}(a_{b},r)$ when possible as long as $r>{\rm max}(a_{b},20r_{g}).$
We initialized each disk with a surface density profile of $\Sigma=\Sigma_{0}\exp[-(r/2.5a_{0})^{-12}][1-0.7\sqrt{(}a_{0}/r)]$ (see, e.g. Lynden-Bell and Pringle, 1974; Miranda et al., 2017; Dittmann and Ryan, 2022), a radial velocity of $3\nu/2r$, and an angular velocity profile of
@@ IV Results
For the sake of brevity we do not report here a detailed analysis of gas torques on the binary during the inspiral. However, over the course of decoupling gravitational torques between the binary, circumbinary disk, and streams of gas within the cavity, which are largely responsible for negative contributions to the gravitational torque on the binary from the gas (e.g., Tiede et al., 2020; Dittmann and Ryan, 2022), become weaker as the binary decouples. The minidisks and their positive torque contributions remain, so the gravitational torque on the binary tends to increase by a factor of order unity (normalized by the accretion rate and binary angular momentum $\dot{M}J_{b}/M$) during the decoupling process, although this effect may be more substantial in thinner disks.
IV.1 Disk Morphology
@@
We measured the evolution of the circumbinary disk cavity using time-averaged histograms of the fluid surface density, binned according to fluid element semi-major axis (see, e.g. Figure 17 of Dittmann and Ryan, 2022). We defined the semi-major axis of the cavity ($a_{\rm cav}$) as that at which surface density reaches a value of $\Sigma_{0}/5$ at $a>a_{b}$. Although our definition of cavity semi-major axis is somewhat arbitrary, we have verified that our results are insensitive to the precise definition used. The top panel of Figure 2 presents the results of these measurements for each viscosity, along with nonlinear least-squares fits of a smoothly broken power law shown using dashed lines. For our broken power law fit, we fix early- and late-time power law indices $a_{\rm cav}\propto a_{b}^{p}$ to $p=1$ and $p=0$ respectively, fitting for a break location, transition width, and constant of proportionality. We define the binary semi-major axis which characterizes the decoupling of the binary from the circumbinary disk ($a_{d}$) as the location of the power-law break. Although $a_{d}$ is a characteristic point in the evolution of the binary, Figure 2 also illustrates that decoupling is a gradual process as opposed to an instantaneous one.
The bottom panel of Figure 2 compares our measurements with the predictions made by Equations (5) and (6) using $\lambda=1$ and measuring $\xi$ at the end of the burn-in stage of each simulation. As expected based on Section II.1, timescale-based reasoning overestimates the characteristic binary semi-major axis at which decoupling occurs. Additionally, Equation (6) makes fewer and weaker assumptions than Equation (5), not depending on $\lambda$ and depending more weakly on $\xi$. Comparing the values of $a_{d}$ in Figure 2 to the surface density snapshots in Figure 1 illustrates that the disk-binary system continues to evolve together after this nominal decoupling point ($a_{d}\approx 56\,R_{g}$ for the $\nu_{0}=0.003$ simulation), albeit at a slower rate.
IV.2 Accretion
@@
In the $\nu_{0}=0.001$ case, periodicity is dominated by the orbital period of the disk cavity and its harmonics as has been observed previously (e.g. Farris et al., 2014; Noble et al., 2021). Although in the lab frame the cavity orbital frequency slightly increases between the beginning of the simulation and decoupling, as the cavity shrinks, the frequency relative to the binary decreases, and the binary evolution outpaces that of the disk. At late times, almost all lab-frame variability is dwarfed by the secular decline in the accretion rate. Our averaging window was also large enough, in terms of the binary orbital period, to identify some low-frequency periodicity due to accretion from the cavity walls. However, even in terms of the evolving binary period, there is virtually no variability on the orbital period of the binary. In these cases, however, effects such as Doppler modulation (e.g., D’Orazio and Haiman, 2017) and self-lensing (e.g., Davelaar and Haiman, 2022) might lead to variability in observed light curves.
IV.3 Gravitational Waves